The measure of an angle is 1/3 of its supplement.find the measure of the angle.

Answer by stanbon(75887)

The measure of an angle is 1/3 of its supplement.find the measure of the angle
Let the measure of the angle be "x".
It's supplement is 180-x.
----
Equation:
x = (1/3)(180-x)
x = 60 - (1/3)x
(4/3)x = 60
(1/3)x = 15
x = 45
--------------
Cheers,
Stan H.
==============